examine the f:r->r defined by f(x)={1/6(x+1)^3 x is not equal to 0 5/6 x=0} for continuity on R.If it is not continuous at any point of R,find the nature of discontinuity there

1
Expert's answer
2022-01-10T19:12:17-0500
"f(x)=\begin{cases}
 \dfrac{1}{6}(x+1)^3 , x\not=0\\
\\
 \dfrac{5}{6}, x=0
\end{cases}"

The function "f(x)" is continuous on "(-\infin, 0)\cup (0, \infin)" as polynomial.

"\lim\limits_{x\to0}f(x)=\lim\limits_{x\to0}\dfrac{1}{6}(x+1)^3=\dfrac{1}{6}"

"\lim\limits_{x\to0}f(x)=\dfrac{1}{6}\not=\dfrac{5}{6}=f(0)"

The function "f(x)" has a removable discontinuity at "x=0."